# Language Models are Crossword Solvers

### v2

@@ 3.4 Sub-token Counting
Despite massive performance gains demonstrated by current SoTA LLMs, one area in which they perform poorly is adherence to length constraints, suggesting they lack the ability to perform sub-token counting i.e. counting the number of characters present in a word or phrase. We observed that even the best performing model (GPT-4-Turbo) produces answers of incorrect length on 26.2% and 16.9% of the Init and NYT datasets respectively.
The fact that LLMs are stumped by this rather simple task is on the surface a little baffling, however, we must consider that most SoTA LLMs first tokenize sequences using some Byte-Pair Encoding (Sennrich et al., 2016) like scheme. Thus, during the word embedding stage in Transformers (Vaswani et al., 2017), when the token is converted to an embedding vector, information about constituent characters in the token is lost, and must be regained somehow during training. Although we are not sure how LLMs regain this length information during training, we have a hunch.
@@
There are websites  [^5] https://word.tips/words-by-length/ for example. which contain large lists of words with their corresponding lengths. Often replies in message boards also include a count of the number of characters in the reply. Artifacts like these, which contain enough information to infer length of tokens, go on to become part of the datasets that LLMs are trained on. We hypothesise that LLMs learn to count sub-tokens based on this information provided during training.
@@
To investigate this further, we devised the sub-token counting task, wherein the LLM is provided a sequence of (lower case) characters without white-spaces and asked to predict the number of characters making up the sequence. To test our hypothesis, we consider three sets of 1000 (English) words - Common, Medium and Rare based on word unigram frequencies curated by Segaran and Hammerbacher (2009) from Google’s Trillion Words corpus. The Common, Medium and Rare words have ranks in the range of 1-5,000, 47,500-52,500 and 95,000-100,000 respectively.
@@
Figure 3: LLMs ability to count the number of characters in a word decline with frequency of the word.
@@
If our hypothesis is incorrect, and language models have a widely generalizable ability to perform sub-token counting, we should see no difference in counting performance across words with different prevalence. However, we observe that (see Figure 3) the accuracy of LLMs at the sub-token counting task declines with frequency of the token, for all LLMs tested. We further analyzed if there was a difference in sub-token counting performance between words that are part of the model vocabulary and randomly generated gibberish following the same distribution of lengths, to account for potential distribution shifts between frequent and rare words. To do this we first created a set of words by taking an intersection of all words that are part of every open source model in consideration and the list of top 100,000 words. This is to ensure that the sequences are - (a) extremely likely [^6] We can’t be sure for the proprietary models to be vocabulary tokens for every model in consideration and (b) are not special tokens like <bos>. Then we created a set of gibberish words by replacing each character of the vocab set words with a randomly chosen character from the set {a-z}, thus guaranteeing that they have the same length distributions.
@@
Table 2: LLM sub-token counting performance for vocabulary words and gibberish. In addition to declining counting performance with decreasing prevalence of the word, there is a sizeable difference in counting performance between vocabulary words and randomly generated gibberish.
We find (see Table 2) that in addition to counting accuracy being affected by frequency, there is often a huge disparity between the accuracy of vocabulary vs gibberish words. Although, this does not conclusively show that LLMs rely on memorized training instances to perform sub-token counting, it does provides strong evidence suggesting that LLMs learn to count based on length information containing artifacts in training data. Future works exploring this idea further would be compelling.
@@ 4 SweepClip - Our proposed algorithm to solve crosswords with LLMs
Emboldened by our successes in deciphering answers from crossword clues with LLMs, we set out to address the challenge of tackling the problem of filling crossword grids with LLM assistance. Note that this task not only involves generating correct answers from provided clues, but hinges on exploiting constraints from already generated words, and back-tracking to eliminate past generations that do not fit well when new evidence becomes available. Since LLMs demonstrate this ability, when paired with the right search algorithm, it should be possible to solve to crosswords with the aid of LLMs.
Algorithm 1 - SweepClip (Sketch)
1:Given crossword grid, clues and an LLM.
2:Generate candidate answers for all clues. (sweep)
[…]
4:     Prune answers not in largest connected component. (clip)
5:     Prune conflicting answers.
6:     Generate answers for neighbors of surviving answers.
7:     if solved or budget_exceeded then
8:         break
9:     end if
10:end for
Our proposed algorithm (outlined in Algorithm 1 and detailed in Appendix A) first generates a set of candidate answers for all clues provided with the crossword (sweep) and uses a graph based criteria to eliminate answers that do not fit well to the grid (clip). Following this, we use the constraints generated from the previous step answers, generate more candidate answers, and prune the bad fitting candidates. We iteratively apply this strategy, until either - (i) the entire crossword is filled, (ii) we exceed a preset number of iterations, or (iii) we run out of LLM computational budget.
For pruning we use the largest connected component from the answers generated so far to ensure that the generated answers are somewhat coherent amongst themselves. This is over-restrictive, as it is possible that isolated answers are correct, however, we find that this strict pruning strategy works better to eliminate bad answers early rather than using the constraints imposed by potentially bad answers to generate further bad answers.
We would like to highlight that our algorithm is sub-optimal, and does not provide convergence or correctness guarantees beyond elimination of conflicting answers. Note that this problem is NP-Hard and it is difficult to find approximate solutions for it (Kulshreshtha et al., 2022). This algorithm was designed from the ground up to minimize cost incurred from calls to the LLM by avoiding branching based on LLM answers.
@@ 5 Results
In this section we present results from our algorithm at the crossword solving task and present further results investigating the performance demonstrated by current generation SoTA LLMs.
@@ 5.1 Solving Cryptic Crossword Clues
We observed a massive performance increase in cryptic crossword clue deciphering in the current generation of SoTA LLMs. We present comparisons between the results we obtained on the cryptic crossword clue answering task 777treated as QA with the Init dataset, and previously reported SoTA results in Table 3.
@@
Table 3: Comparison of our results with previously reported SoTA Results.
We also note that Chain-of-thought (Wei et al., 2022b) prompting (1 shot) with self-consistency Wang et al. (2023) (3 samples) leads to further performance gains. Our best result on the word-init-disjoint split is 20.85% which improves over the previous SoTA (9.5%) by a factor of 2.2$\times$ without any fine-tuning.
5.2 Solving Crossword Puzzles
For this task we employed our algorithm SweepClip on a set of 100 randomly sampled Monday NYT crossword puzzles. We used two LLMs for this - GPT-4-Turbo with max_depth of 30 and a budget of 0.5 USD per crossword, and LLaMA 3 70B with max_depth of 35 and a budget of 600 LLM calls.
Error Tolerance % of Crosswords
LLaMa 3 GPT-4 T
100% solved 0 48
$\leq$ 1 character error 1 55
$\leq$ 5 character error 10 71
$\geq 90\%$ Accuracy 30 80
$\geq 50\%$ Accuracy 82 98
We find (see Table 4) that our algorithm with GPT-4-Turbo solves 48% of crosswords without any errors, and 55% of crosswords with at most 1 character wrong. The average character level accuracy in crossword solving is 93.1% ($\pm$ 14.1%). Our algorithm improves the clue-wise answer accuracy 888clue level accuracy is different from character level accuracy (Kulshreshtha et al., 2022), e.g. it is possible to have a filled in crossword without deciphering all clues. (exact match) to 89.6% ($\pm$ 16.9%) from the base accuracy (without the algorithm) which is 43.5% ($\pm$ 23.5%) an improvement of 2.1$\times$. Note the previously reported SoTA accuracy on this task was 26% with retrieval augmented generation and an SMT solver coupled with an oracle that eliminates parts of the crossword grid that do not have suitable generated answers (Kulshreshtha et al., 2022).
The performance for the smaller LLaMA 3 70B is worse, however, our algorithm still manages to improve final clue answering accuracy to 59.4% ($\pm$ 24.1%) from a base accuracy of only 22.3% ($\pm$ 14.4%).
Our algorithm implicitly performs self-consistency checks to improve candidate answers. For example the generated clue at the first sweep may be discarded and regenerated with a different prompt and then accepted if it does not produce conflicts, and is coherent with a larger number of other answers. Thus with the application of our algorithm, we have successfully exploited the constraint information to boost the performance of LLMs far beyond what would have been possible with straightforward QA like clue deciphering. To the best of our knowledge this is the first algorithm that demonstrates succesfully solving crosswords with the aid of LLMs.
@@ 5.3 Data Contamination and Generalizability
Having seen remarkable performance improvements in the current generation of SoTA LLMs, we set out to investigate the reason behind this success. To rule out data-contamination, we curated additional cryptic crossword clue datasets comprising entirely of puzzles published after May 20, 2024, which is after the knowledge cutoff date of all LLMs examined. These datasets are sourced from The Guardian [^9] https://www.theguardian.com/ and Lovatts Puzzles [^10] https://lovattspuzzles.com. Note, the Init dataset is also sourced from the The Guardian, thus the results reported in Table 5 should be consistent.  [^11] More details on data curation is in Appendix B.
@@
Table 5: Performance of LLMs on curated datasets that appeared after knowledge cut-off. Note, Init by Rozner et al. (2021), also sourced their data from The Guardian, thus these results provide a fair head-to-head comparison of performance.
We see no appreciable difference in performance on the post-cutoff dataset, (see Table 5) leading us to conclude that these LLMs can generalize beyond potential contamination in their training set. The performance of models on the Lovatts Puzzles dataset (post-cutoff) is even better than that on the Guardian dataset.
To ascertain if the models can reason about unseen cryptic crossword clues to arrive at correct responses, we elect to perform human evaluation121212 Further details in Appendix E of model responses. We employ a 3-shot Chain-of-thought prompt to elicit a reasoned response to crossword clues from GPT-4-Turbo, which are then analyzed by our team for soundness vis-à-vis factual and logical errors. We choose 100 samples from the post-cutoff Lovatts set and rely on consensus of all evaluators to ascertain if a response contains factual or logical errors (e.g. wrong counting, wrong anagram, drawing a conclusion that does not follow from the premise).
Sound Unsound
(61%) (39%)
Correct (65%) 48% 17%
Wrong (35%) 13% 22%
Table 6: Results from human evaluation of reasoning provided by GPT-4-Turbo. An answer is called correct if the final answer provided by the model matches the ground truth exactly. The answer is called sound if it contains no logical or factual errors.
@@
6 Conclusions and Discussions
Solving crosswords requires proficiency in a multitude of desirable skills, and cryptic crosswords were previously thought to be firmly in the domain of human experts. Our findings challenge this notion. We have demonstrated that the current generation of SoTA LLMs show significantly improved aptitude at solving straight and cryptic crossword puzzle clues, and can exploit constraints provided by partially solved crosswords to boost this performance further. We also found that this emergent ability generalizes well to the post knowledge cut-off regime and is accompanied with the capacity to produce reasoned explanations.
We have also developed an algorithm that, with the aid of LLMs, further boosts this performance, and achieves 93% accuracy on Monday New York Times crosswords. With the aid of this algorithm it is possible to achieve massive performance gains even if the baseline LLM accuracy is low, thus indicating that, when paired with the right search strategy, LLMs can succesfully solve crosswords.
This calls into question previously reported results which claim that LLMs massively lag behind human experts at solving cryptic crosswords. We are firmly of the opinion that the current gap between human experts and LLMs will cease to exist in the near future. Incorporating length constraints effectively still remains a challenge, and further enhancements in numerical reasoning abilities in LLMs is a promising avenue to pursue.
@@ 6 Conclusions and Discussions
Limitation 1 - Reporting crossword solving results We only reported results on solving Monday New York Times crossword puzzles and didn’t report Tuesday-Thursday which typically increase in level of difficulty. We also did not solve cryptic crossword grids. There are primarily two reasons for this. (i) As we have seen in the paper, even if the base LLM has relatively low accuracy when paired with a search/prune algorithm, the accuracy can be improved considerably. There is a trade-off between base accuracy and the number of LLM calls, i.e. if the base accuracy is lower, the algorithm needs to make a lot more calls to the LLM increasing computational costs. We have observed that it is possible to solve harder NYT (straight) crosswords or cryptic crosswords, but at an increased computational budget. This makes reporting statistics cost-prohibitive for us at this time. (ii) The cryptic crossword datasets do not come with grid information, and instead provide cryptic clues and answers as question-answer pairs. To successfully demonstrate cryptic crossword grid solving we would need to acquire a high quality cryptic crossword dataset with full grid information.
@@
Limitation 2 - Algorithm Our algorithm is potentially sub-optimal as it discards potential correct answers and does not fully explore the consequences of each possible generated answer. This is again a computational cost saving technique. Ideally, we would generate one candidate answer, followed by all its neighbors, and whenever there is a conflict, branch out and explore all options to see which is a better fit. We elect not to do this, because it involves a potentially exponential number of LLM calls. Note that this problem is NP-Hard, however, future studies with much larger computational budget for LLM calls will definitely have more success employing a more thorough search strategy.
@@
Limitation 3 - Sub-token Counting We have demonstrated that LLMs show limitations at sub-token counting, and have hypothesised that they do not have widely generalizable skill at this task, rather relying on memorization. The evidence provided in this paper is compelling, but definitely not enough to conclusively prove this hypothesis. Further studies are required that can intervene at the pre-training stage of LLMs to conclusively demonstrate whether this hypothesis is True. We do not possess the computational power to undertake such a study at this stage.
@@
Limitation 4 - Fine-tuning We do not report results with fine-tuned models. It is possible that further improvements could be seen with fine-tuning, however, we are interested in studying the emergent properties of large language models in this context instead of building special purpose models that can solve crosswords.
@@
In keeping with ACL ethical guidelines we make all scientific artifacts generated for this study freely available and open source under the MIT licence, this includes all software we created, every prompt that was used, all raw model outputs, and data we generated. We do not distribute the New York Times crossword dataset since it is the intellectual property of The New York Times. We have purchased a subscription to access their puzzle repository, but used the data without their consent for this study. The use of their data falls within the terms of the Fair Use doctrine set within 17 U.S.C. §107. To aid in reproducibility we provide a list of dates to uniquely identify the crossword puzzles used in the study, and provide detailed instructions on how to acquire them through the New York Times.

### v3

@@ 3.4 Sub-token Counting
Despite significant performance gains, SoTA LLMs struggle with adherence to length constraints, suggesting an inability to count characters within words or phrases (sub-token counting). We observed that even the best performing model, GPT-4-Turbo, produces answers of incorrect length on 26.2% of the Init dataset and 16.9% of NYT dataset. This may be explained by the tokenization methods used in LLMs, such as Byte-Pair Encoding (Sennrich et al., 2016). During the word embedding stage in transformers (Vaswani et al., 2017), tokens are converted into embedding vectors, causing the loss of information about individual characters. This character-level information must be relearned during training. While we are unsure exactly how LLMs regain this information, we suspect they learn from training data that include explicit length details.
@@
There are websites [^5] https://word.tips/words-by-length/ for example. that contain large lists of words with their corresponding lengths. Often replies in message boards also include a count of the number of characters in the reply. Artifacts like these, which contain enough information to infer the length of tokens, go on to become part of the datasets that LLMs are trained on. We hypothesize that LLMs learn to count sub-tokens based on this information provided during training.
@@
To investigate this further, we devised the sub-token counting task, wherein the LLM is provided a sequence of (lowercase) characters without whitespaces and asked to predict the number of characters making up the sequence. To test our hypothesis, we consider three sets of 1000 (English) words–Common, Medium and Rare–based on word unigram frequencies curated by Segaran and Hammerbacher (2009) from Google’s Trillion Words corpus. The Common, Medium and Rare words have ranks in the range of 1-5,000, 47,500-52,500, and 95,000-100,000, respectively.
@@
Figure 3: LLMs ability to count the number of characters in a word declines with the frequency of the word.
@@
If language models have a widely generalizable ability to perform sub-token counting, we should see no difference in counting performance across words with different prevalence. However, we observe that (see Figure 3) the accuracy of LLMs at the sub-token counting task declines with the frequency of the token for all LLMs tested. We further analyzed if there was a difference in sub-token counting performance between words that are part of the model vocabulary and randomly generated gibberish following the same distribution of lengths to account for potential shifts of distribution of length between frequent and rare words. To do this, we first created a set of words by taking an intersection of all words that are part of the vocabulary of every open-source model in consideration and the list of the top 100,000 words. This is to ensure that the sequences are extremely likely [^6] We can’t be sure about the proprietary models. to be vocabulary tokens for every model in consideration and are not special tokens like <bos>. Then we created a set of gibberish words by replacing each character of the vocab set words with a randomly chosen character from the set {a-z}, thus guaranteeing that they have the same length distributions.
@@
Table 2: LLM sub-token counting performance for vocabulary words and gibberish.
We find (see Table 2) that in addition to counting accuracy being affected by frequency, there is often a large disparity between the accuracy of vocabulary vs. gibberish words. Although this does not conclusively show that LLMs rely on memorized training instances to perform sub-token counting, it does provide strong evidence suggesting that LLMs learn to count based on length information containing artifacts in training data. Future works exploring this idea further would be compelling.
@@ 4 SweepClip - Our proposed algorithm to solve crosswords with LLMs
In this section, we address the problem of filling crossword grids with LLM assistance. Note that this task not only involves generating correct answers from provided clues but also hinges on exploiting constraints from already generated words and backtracking to eliminate past generations that do not fit well when new evidence becomes available. Since LLMs demonstrate this ability, when paired with the right search algorithm, it should be possible to solve crosswords with the aid of LLMs.
Our proposed algorithm (detailed in Appendix A) first generates a set of candidate answers for all clues provided with the crossword (sweep) and uses a graph-based criterion to eliminate answers that do not fit (clip). Then we use the constraints generated from the previous step answers to generate more candidate answers777with appropriate masks provided as prompts, see Appendix A. and prune the bad-fitting candidates. We iteratively apply this strategy until either (i) the entire crossword is filled, (ii) we exceed a preset number of iterations, or (iii) we run out of LLM computational budget.
For pruning, we use the largest connected component from the answers generated so far to ensure that the generated answers are somewhat coherent amongst themselves. This is over-restrictive, as it is possible that isolated answers are correct; however, we find that this strict pruning strategy works better to eliminate bad answers early rather than using the constraints imposed by potentially bad answers to generate further bad answers (see Appendix A).
@@ 5 Results
In this section, we present results from our algorithm at the crossword-solving task and present further results investigating the performance demonstrated by current-generation SoTA LLMs.
5.1 SweepClip–Solving NYT Crossword Puzzles
For this task, we employed our algorithm, SweepClip, on a set of 100 randomly sampled Monday NYT crossword puzzles. We used two LLMs for this: GPT-4-Turbo and Llama 3 70B (see Appendix A).
Error Tolerance % of Crosswords
LLaMa 3 GPT-4 T
100% solved 0 48
$\leq$ 1 character error 1 55
$\leq$ 5 character error 10 71
$\geq 90\%$ Accuracy 30 80
$\geq 50\%$ Accuracy 82 98
We find (see Table 3) that our algorithm with GPT-4-Turbo solves 48% of crosswords without any errors and 55% of crosswords with at most 1 character wrong. The average character level accuracy in crossword solving is 93.1% ($\pm$ 14.1%). Our algorithm improves the clue-wise answer accuracy888clue-level accuracy is different from character-level accuracy (Kulshreshtha et al., 2022); e.g., it is possible to have a filled-in crossword without deciphering all clues. (exact match) to 89.6% ($\pm$ 16.9%) from the base accuracy (without the algorithm), which is 43.5% ($\pm$ 23.5%), an improvement of 2.1$\times$. Note, the previously reported SoTA accuracy on this task with a foundational LM (without fine-tuning) was 26% with retrieval-augmented generation and an SMT solver coupled with an oracle that eliminates parts of the crossword grid that do not have suitable generated answers (Kulshreshtha et al., 2022).
The performance for the smaller LLaMA 3 70B is worse; however, our algorithm still manages to improve final clue answering accuracy to 59.4% ($\pm$ 24.1%) from a base accuracy of only 22.3% ($\pm$ 14.4%). Note that this result, alongside those presented in Section 3, serve to ablate our crossword-solving approach.
Thus, with the application of our algorithm, we have successfully exploited the constraint information to boost the performance of LLMs beyond what would have been possible with straightforward QA like clue deciphering. To the best of our knowledge, this is the first algorithm that demonstrates successfully solving crosswords with the aid of an out-of-the-box LLM.
@@ 5.2 Solving Cryptic Crossword Clues
Table 4: Comparison of our results with previously reported SoTA results. Results are on the Init dataset with crossword clue deciphering treated as QA.
We observed (see Table 4) that SoTA LLMs have significantly improved cryptic crossword clue deciphering abilities. We also note that chain-of-thought (Wei et al., 2022b) prompting (1-shot) with self-consistency Wang et al. (2023) (3 samples) leads to further performance gains. Our best result on the word-init-disjoint split is 20.85%, which improves over the previous SoTA (9.5%) by a factor of 2.2$\times$ without any fine-tuning.
@@ 5.3 Data Contamination and Generalizability
To rule out data contamination as the reason behind the significant performance gains by SoTA LLMs, we curated additional cryptic crossword clue datasets comprising entirely of puzzles published after May 20, 2024, which is after the knowledge cutoff date of all LLMs examined. These datasets are sourced from The Guardian and Lovatts Puzzles. The answers in these post-cutoff sets were checked against the combination of all other cryptic crossword datasets employed in the study (665,497 answers in total) to check for potential duplicates. None was found for the post-cutoff Guardian set, and 2 were found for the post-cutoff Lovatts set, which were removed. Note: the Init dataset is also sourced from The Guardian; thus, the results reported in Table 5 should be consistent [^9] More details on data curation is in Appendix B..
@@
Table 5: Performance (accuracy) of LLMs on curated datasets that appeared after the knowledge cut-off. Note, Init by Rozner et al. (2021), also sourced their data from The Guardian; thus, these results provide a fair head-to-head comparison of performance.
We see no appreciable difference in performance on the post-cutoff dataset (see Table 5), leading us to suggest that these LLMs can generalize beyond potential contamination in their training set.
5.4 Human Evaluation and Further Analysis
To ascertain if the models can reason about cryptic crossword clues to arrive at correct responses, we elect to perform human evaluation101010 Further details in Appendix E of model responses. We employ a 3-shot chain-of-thought prompt to elicit a reasoned response to crossword clues from GPT-4-Turbo, which are then analyzed by our team for soundness vis-à-vis factual and logical errors. We choose 100 samples from the post-cutoff Lovatts set and rely on the consensus of all evaluators to ascertain if a response contains factual or logical errors (e.g., wrong counting, wrong anagram, drawing a conclusion that does not follow from the premise, etc.). The author-annotated reasoning responses had low inter-observer variability (Fleis’ $\kappa=0.94$, pre-consensus) and have been made publicly available.
GPT 4 Turbo Sound $\neg$¬Sound
(61) (39)
Correct (65) 48% 17%
Wrong (35) 13% 22%
Table 6: Results from human evaluation of reasoning. Chain-of-thought elicited reasoning responses from GPT-4-Turbo are evaluated for soundness. An answer is called correct if the model prediction exactly matches the ground truth. The answer is called sound if it contains no logical or factual errors. Results are on the post-cutoff Lovatts’ set.
GPT-4-Turbo ANG SCJ CNT HOM OTH
(27) (26) (18) (8) (21)
EM 74% 50% 56 % 100 % 67 %
Model ANG CNT
(8.5%) (2.5%)
Llama 3 2.6% 12.2%
Claude 3 7.9% 15.3%
GPT 4 T 25.0% 33.7%
Table 7: Do LLMs have common “failure modes”? ANG refers to anagrams, SCJ refers to synonym conjugation, CNT refers to containment, HOM refers to homophones, and OTH refers to others. (Left) shows results from GPT-4-Turbo on the post-cutoff Lovatts set, and (right) shows results on the combined Init and Cryptonite datasets. The numbers in parentheses refer to prevalence of the clue type.
@@ 5.4 Human Evaluation and Further Analysis
Furthermore, to analyze if LLMs demonstrate any common failure modes, we manually annotated the human evaluation dataset based on the principal skill required to solve a particular puzzle clue (see Table 7). ANG indicates that the answer is an anagram of some words of the clue, HOM indicates that the answer is a homophone of some words of the clue, CNT indicates that the answer is disguised in a contiguous section of the clue, and SCJ indicates that the answer is found by combining several words that are synonyms of various parts of the clue. The class OTH lumps together a variety of other kinds of skills (e.g., spoonerisms, acronyms, world knowledge, and various other kinds of character manipulations). Examples of each are provided in Appendix E.
Owing to the limited number of human annotations, the results (Table 7 left) are not statistically significant, but the trends suggest that GPT-4-Turbo may exhibit strong performance on anagrams (74% accuracy vs. 65% baseline [^11] Baseline refers to mean accuracy across all kinds of clues.) and homophone-based clues (100% accuracy). We also attempted to perform the same analysis on the Cryptonite and Init datasets (4000 clue-answer pairs) and found similar trends for anagram performance by GPT-4-Turbo (25% accuracy vs. 21.1% baseline). However, this does not hold for Llama 3 70B (2.6% accuracy vs. 7.16% baseline) and Claude 3 Sonnet (7.9% accuracy vs. 11.85% baseline), suggesting that Llama 3 70B and Claude 3 Sonnet have a diminished ability to deal with anagrams. Note that unlike (ANG, CNT), (HOM, SCJ, OTH) cannot be automatically detected reliably, which limits broader analysis for these kinds of clues.
To quantify the effect of sub-token counting performance on clue-solving performance, we devise a further experiment. We consider all such clues for which the model correctly deduced the semantics of the clue but failed to adhere to the length constraints (e.g., LECTURER instead of PROFESSOR or NANNA instead of GRANNY). We consider wrong LLM predictions with a semantic similarity of 0.5 or more  [^12] As given by OpenAI text-embedding-3-large model. and report the percentage of answers of incorrect length. GPT-4-Turbo and Llama 3 70B produce predictions with length errors 46.4% and 59.9% of the time respectively. This suggests that the lack of adherence to length constraints is a major impediment in clue solving for LLMs.
6 Conclusions
Solving crosswords requires proficiency in a multitude of desirable skills, and cryptic crosswords were previously thought to be firmly in the domain of human experts. Our findings challenge this notion. We have demonstrated that the current generation of SoTA LLMs shows significantly improved aptitude at solving straight and cryptic crossword puzzle clues and can exploit constraints provided by partially solved crosswords to boost this performance further. We also found that this emergent ability generalizes well to the post knowledge cut-off regime and is accompanied with the capacity to produce reasoned explanations.
We have also developed an algorithm that, with the aid of LLMs, further boosts this performance and achieves 93% accuracy on Monday New York Times crosswords. With the aid of this algorithm, it is possible to achieve significant performance gains even if the baseline LLM accuracy is low, thus indicating that, when paired with the right search strategy, LLMs can successfully solve crosswords. Such an LLM-guided search approach may be readily adapted to other scenarios requiring certain kinds of constraint satisfaction.
Incorporating length constraints effectively still remains a challenge. We showed that SoTA LLMs struggle with sub-token counting and also that they often provide incorrect predictions despite being semantically close, owing to length constraint violations. This weakness also manifests in other tasks like anagrams, character manipulations, etc., which are heavily reliant on arithmetic abilities. Future research attempting to address this shortcoming would be compelling.
@@ 6 Conclusions
Limitation 1 (Algorithm) - We would like to highlight that our algorithm is sub-optimal, as it discards potential correct answers and does not fully explore the consequences of each possible generated answer. Ideally, we would generate one candidate answer, followed by all its neighbors, and whenever there is a conflict, branch out and explore all options to see which is a better fit. We elect not to do this, because it involves a potentially exponential number of LLM calls. Our algorithm does not provide guarantees of convergence or correctness beyond elimination of conflicting answers. Note that this problem is NP-Hard and it is difficult to find approximate solutions for it (Kulshreshtha et al., 2022). This algorithm was designed to minimize computational cost incurred in LLM calls; however, future studies with much larger computational budget for LLM calls will definitely have greater success employing a more thorough search strategy.
Limitation 2 (Cryptic Crossword Solving) - Compared to straight crosswords, the baseline accuracy in clue answering by LLMs for cryptic crosswords is much lower (e.g.,  18% for GPT-4-Turbo on cryptics vs  41% on straight). Thus, solving cryptic crosswords with current SoTA LLMs would require a much more extensive search strategy. Our current approach with a budget limit of 0.5 USD per crossword is unsuccessful (12% of cryptic crosswords solved with 50% letter accuracy) at solving cryptics, and our financial constraints do not permit a more thorough investigation with improved algorithms at this time.
Limitation 3 (Reporting Crossword Solving Results) - We only reported results on solving Monday New York Times crossword puzzles and did not report Tuesday-Thursday which typically have a higher level of difficulty. There are primarily two reasons for this. (i) As we have seen in the paper, even if the base LLM has relatively low accuracy, when paired with a search/prune algorithm, the accuracy can be improved considerably. There is a trade-off between base accuracy and the number of LLM calls, i.e., if the base accuracy is lower, the algorithm needs to make a lot more calls to the LLM increasing computational costs. We have observed that it is possible to solve harder NYT (straight) crosswords, but at an increased computational budget. This makes reporting statistics cost-prohibitive for us at this time.
@@
Limitation 4 (Sub-token Counting) - We have demonstrated that LLMs show limitations at sub-token counting, and have hypothesised that they do not have widely generalizable skill at this task, rather relying on memorization. The evidence provided in this paper is compelling, but definitely not enough to conclusively prove this hypothesis. Further studies are required that can intervene at the pre-training stage of LLMs to conclusively demonstrate whether this hypothesis is true. We do not possess the computational power to undertake such a study at this stage.
@@
Limitation 5 (Fine-tuning) - We do not report results with fine-tuned models. It is possible that further improvements could be seen with fine-tuning, however, we are interested in studying the emergent properties of large language models in this context instead of building special purpose models that can solve crosswords.
@@
In keeping with ACL ethical guidelines we make all scientific artifacts generated for this study freely available and open source under the MIT licence, this includes all software we created, every prompt that was used, all raw model outputs, and data we generated. The test sets used for all experiments are also made availabe, with the exception of the full grids of the New York Times crossword puzzles. We do not re-distribute the (publicly available) New York Times crossword dataset since it is the intellectual property of The New York Times. The use of their data falls within the terms of the Fair Use doctrine set within 17 U.S.C. §107. To aid in reproducibility we provide a list of dates to uniquely identify the crossword puzzles used in the study, and provide detailed instructions on how to acquire them through the New York Times. We do not use any data to train any model in this study.
@@
Acknowledgments
The authors would like to thank Akshay Chaturvedi and Joy Mahapatra for their helpful feedback on this paper. This research was funded in part by the Indo-French Centre for the Promotion of Advanced Research (IFCPAR/CEFIPRA) through project number CSRP 6702-2.